OpenStudy (anonymous):

ind an equation of the line that satisfies the given conditions. Through (−1, −3) and (7, 5)

OpenStudy (anonymous):

use the same approach : y1-y=m(x1-x) find slope (m) & write equation

OpenStudy (anonymous):

5-(-3)=m(7-(-1)) 8=m*8 m=1 y-5=1*(x-7) y=x-7+5 y=x-2 or standard form: x-y-2=0

OpenStudy (anonymous):

so x-y-2=0 is the answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

5-10-2=0

OpenStudy (anonymous):

5x+10-2=0?

OpenStudy (anonymous):

-2z

OpenStudy (anonymous):

both are answers: y=x-2 (slope form) or standard form: x-y-2=0 it's the same thing just answer presented differently; one y=... and another expression=0 good?

OpenStudy (anonymous):

but do i plug in the x and y intercept for x and y

OpenStudy (anonymous):

you asked: "find an equation of the line ..." x-y-2=0 (or y=x-2) is equation of the line. the two points were given to help to identify the line - line goes through those points

OpenStudy (anonymous):

if there are no points given - there are infinite amount of lines.. right? so, two points help you to find THE line

OpenStudy (anonymous):

over my head.

OpenStudy (anonymous):

i think the answer is 10x+5y-2=0

OpenStudy (anonymous):

OK. look at the peace of paper - it's your 2 dimensional field. 1)how many lines you can draw on a peace of paper? 2)how many lines your can draw through 2 points?

OpenStudy (anonymous):

10x+5y-2=0 can not be the answer! put the values of x & y (your given points) in your equation -> your statement will be wrong!

OpenStudy (anonymous):

for example point(-1,-3) - first given point means that x=-1 and y=-3 put this values on your equation: 10(-1)+5*(-3)-2=0 -10-15-2=0 -27=0 NOT TRUE statement so, your equation is wrong one. please see my solution - I put it in details for you

OpenStudy (anonymous):

o gotcha thanks